# Improving the Quality of Web-mined Parallel Corpora of Low-Resource Languages using Debiasing Heuristics

### v2

@@ 3 Methodology
Improved Taxonomy for Noise Although translation pairs can have overlapping URLs, acronyms, etc, excessive inclusions of such content in a sentence (e.g. consider the sentence ‘Contact: Diane Anderson 076-8268914, info@sandnasbadenscamping.se’ [^2] More examples are in Table 8 in Appendix A) do not provide meaningful content for an NMT system to learn from. However, under  Ranathunga et al. (2024)’s taxonomy, such sentence pairs would likely be categorized as correct translations. Therefore, we define a new noise category CCN to capture such sentence pairs.
Secondly, we consider the upper limit for short sentences as five words [^3] We consider thresholds 3,4 and 5 and empirically select this threshold as the optimal., based on common observations. Finally, we improve the definition of WL (wrong language) to consider a threshold in determining whether a sentence pair should be marked as wrong language. Table 3 shows these changes. The complete list of these noise categories and example parallel sentences is available in Table 9 and Table 10 (respectively) in Appendix A.
Selection of Heuristics
Table 4 shows how the heuristics discussed in Section 2.3 may help in removing different noise categories. Note that de-duplication based heuristic cannot be associated with any noise category, as it does not apply to individual sentence pairs. In addition to de-duplication strategies discussed in Section 2.3, we introduce an n-gram based de-duplication, meaning that sentences would be removed if they overlap in a consecutive n-gram sequence.
@@
Overlapping Numbers/acronyms/URLs/email CCN LID, sentWRatio, sentCRatio
@@
We conduct a human evaluation to quantify the noise before and after applying the heuristics. The annotator selection criteria, resources and training provided, the payment details, etc are described in Appendix F.
From the top 1000 samples in each of the ranked corpora using LASER3, XLM-R and LaBSE, we randomly select 100 parallel sentences for each language pair. We ask each translator to select an annotation category for each sentence pair using the error taxonomy discussed in Section 3. Each sentence pair was annotated by two translators to reduce potential bias caused by the individual translators. Inter-annotator agreement was calculated using the Pearson correlation coefficient, which resulted in En-Si 0.88, En-Ta 0.68 and Si-Ta 0.49.
@@ 4.1 Data
We use the language pairs, En-Si, En-Ta, and Si-Ta in our experiments. Sinhala and Tamil are morphologically rich, low-resource and mid-resourced languages Joshi et al. (2020); Ranathunga and De Silva (2022), respectively. Languages were selected considering the availability of human evaluators. We select CCMatrix and CCAligned as the web-mined corpora. Both these corpora include parallel data for the language pairs considered in the search. More details of these datasets and language pairs are in Appendix B.
For the NMT experiments, we use the dev and devtest subsets from the Flores+ dataset [^4] https://github.com/openlanguagedata/flores as validation and evaluation sets, respectively. Dataset statistics are provided in Table 5.
@@
4.2 Selection of muliPLMs
We select LASER3, XLM-R, and LaBSE for obtaining embeddings for the sentences to determine the semantic similarity. XLM-R, different to others, was trained purely on monolingual data, but has proven to be useful for cross-lingual tasks as well Choi et al. (2021); Conneau et al. (2020). All three models include En, Si, and Ta. Details of the multiPLMs are shown in Appendix C.
4.3 Heuristic-based PDC Experiments
Each heuristic is applied independently to the source (S), target (T), and both sides (ST). In line with the original sentence alignment conducted for CCMatrix and CCAligned, we treat En as the source side. For Si-Ta, Si is considered the source because it is more common for a Si sentence to be translated to Ta Farhath et al. (2018). Finally, for each multiPLM, the retained sentences are ranked in descending order with cosine similarity.
Deduplication: We consider different granularities of de-duplication. ie. identical de-duplication (dedup), de-duplicate by removing numbers only (nums) and removing both numbers and punctuations (punctsNums). Subsequently, we de-duplicate considering different n-gram spans, i.e. 4-grams, 5-grams, 6-grams and 7-grams.
Length-based: We filter short sentences less than five words [^5] We consider sentence length thresholds 3,4 and 5 and empirically decide this value. Results in Table 11 in Appendix D. While some research has suggested removing extremely long sentences (Minh-Cong et al., 2023; Gala et al., 2023), we found that the percentage of longer sentences is lower and that removing them has a negligible effect. Thus, this result is not reported.
LID-based: We use a public LID model [^6] https://github.com/facebookresearch/fairseq/tree/nllb to predict the language of each sentence. The predicted label is then used as a standalone heuristic (LID) and in combination with its associated prediction probability (LIDThresh), with threshold of 0.7777Thresholds below 0.7 reduce NMT results.
Ratio-based: For STRatio, 0.79-1.39, 0.87-1.62 and 0.85-1.57 were selected as thresholds for En-Si, En-Ta and Si-Ta, respectively. These were determined by calculating the mean and the standard deviation obtained for the validation set in a human-crafted En-Si-Ta trilingual dataset (Fernando et al., 2020). Following observations of Hangya and Fraser (2018), 0.6 was selected as the threshold for sentWRatio and  sentCRatio.
4.4 NMT Experiments
First, a Sentencepiece [^8] https://github.com/google/sentencepiece tokenizer with a vocabulary size of 25000 is trained. Then we use the fairseq toolkit  (Ott et al., 2019) to model and train the transformer-based Seq-to-Seq NMT model until convergence. Hyperparameters used in the NMT experiments are shown in Table 12 in Appendix E. The baseline NMT models are trained on the top 100,000 sentence pairs from the ranked corpus. We use ChrF++ Popović (2017) to report NMT results.
5 Results and Analysis
We report the results obtained for the NMT models trained in the forward direction. The results of the experiments are in Table 6. Heuristic-wise best result is summarized in Table 14 in Appendix G. As evident from these tables, as well as from Figure 1, NMT results across different multiPLMs show a great disparity in the baseline NMT scores for En-Si and En-Ta language pairs. In the following sub-sections, we discuss how the use of heuristics is useful in mitigating this disparity and improving overall NMT results.
5.1 Impact of Heuristics on NMT Results
@@
5.1.1 Impact of De-duplication based PDC
We observe that dedup, irrespective of the heuristic-applied side (S/T/ST), outperforms the baseline 89% of the experiments. Overall, de-duplication considering both source and target seems to be the most effective - it outperforms the baseline in 94% of the experiments. In comparison, dedup target only and source only outperform the baseline in 89% and 83% of the experiments, respectively.
We apply our newly introduced ngram-based de-duplication dedup+ngram, on top of  dedup. We find that for each result column, there is a dedup+ngram result that outperforms the best results obtained with the corresponding dedup result. To observe the impact of n value on the NMT result, we plot Figure 2 showing the percentage of ngram experiments exceeding the highest dedup, with respect to the multiPLM (language-wise result is in Figure 4 of the Appendix G). We observe a consistent pattern - n=5 or 6 perform the best in a majority of cases. We believe that 4-gram results in an overly aggressive de-duplication. However, the exact ngram depends on the corpus characteristics.
@@ 5.1.1 Impact of De-duplication based PDC
We observe that dedup+punctNums outperforms  dedup+nums and dedup in 78% and 89% of the experiments (respectively), proving that (dedup+punctNums) to be more impactful. Finally, we analyse the impact of dedup+puntsNums+ngram999 puntsNums and ngram has been applied on top of dedup.. Compared to other dedup combinations101010dedup, dedup+ngram, nums, and dedup+punctNums, dedup+punctNums+ngram produce the best result across 67% of the experiments.
@@ 5.1.2 Impact of Length-Based PDC
sLength surpasses the baseline 89% of the experiments. Therefore, we can conclude that sLength is favourable as a heuristic. When analyzing the side on which the heuristic is applied, we observe that applying it to both ST is the most effective (similar to dedup), followed by T and then S (56%, 28%, and 17% of the experiments, respectively).
Recall that our manual inspection noticed that XLM-R and LaBSE tend to prioritize shorter sentences. This observation is affirmed by the sLength results. For En-Si and En-Ta, this heuristic resulted in substantial gains for many of the XLM-R and LaBSE experiments, while it shows marginal improvements for LASER3.
@@ 5.1.3 Impact of LID-Based PDC
With LID-based PDC, we observe substantial improvements for XLM-R and LaBSE, except for Si-Ta, for which the gains are marginal. Gains are reported by LASER3 as well, though not very significant. The highest gain of +20.61 ChrF++ is reported for XLMR for the CCAligned-EnTa corpus, while a gain of +11.40 is reported by LaBSE for the same corpus.
NMT models trained after applying  LID outperform the baseline in 85% of the experiments, while  LIDThresh outperforms LID in 72% of the experiments. Therefore, we conclude that LIDThresh would be the most suitable heuristic. We observe that the gains for LIDThresh with respect to LID are least with Si-Ta with 50% while for En-Si and En-Ta it is 83% each. We assume this is due to a limitation with the LID model, which is not optimized for Si and Ta.
@@ 5.2 Summary of Heuristic-based PDC
1. Impact of the Individual Heuristics on the NMT Results: The LID-based heuristic emerged as the most impactful in 44% of the experiments. De-duplication and sentence-length heuristics were the most effective in 33% and 17% of the experiments, respectively. The ratio-based heuristic alone did not hold superior results compared to others, making it the least impactful single heuristic. There is no single heuristic that consistently produces the best gains for all scenarios. For LASER3, XLM-R, and LaBSE, we observe the highest ChrF++ gains of +10.57, +20.61, and +11.40 for CCMatrix-EnTa and CCAligned-EnTa and CCAligned-EnTa corpora, respectively.
@@
2. Impact of Combined Heuristics: The combination of heuristics produced the best score, compared to the best-performing individual heuristic, except for the CCMatrix-SiTa. In this, the combination reduced the dataset size by 54% (Table 15 in Appendix G), resulting in only 98k parallel sentences, which could be the reason for the reduced score.
@@
For CCAligned-SiTa (we exclude CCMatrix due to the dataset size limitation), the gains across multiPLMs are marginal. The same pattern holds with LASER3 across En-Ta and En-Si languages, with the exception of CCMatrix-EnTa corpus. Since LASER3 is already trained with OPUS [^11] https://opus.nlpl.eu/ data, it is believed that LASER3 is already optimized Moon et al. (2023) towards ranking the CCMatrix and CCAligned corpora better. This is further affirmed with the relatively higher NMT scores of the baseline experiments.
@@
dedup+punctNums+(n)gram+sLength+LIDThresh combination appears as the best combination in general. The specific n-value is dependent on the dataset. Only in CCAligned-SiTa with XLM-R, the combination performed best without the LID heuristic. However, we include LIDThresh into the combination for two reasons. (1) Even with LIDThreshold, the score only lags by (-0.19) ChrF++ scores, which is negligible. (2) The most influential individual heuristic is LID-based. In the combination, producing the best gains, the ratio-based heuristic was different based on the dataset/language pair. However, in 80% of cases it had been sentWRatio.
3. Reducing the Disparity Across multiPLMs: In order to calculate the disparity between multiPLMs, we calculated the difference between LASER3 and the other two multiPLMs. This is because LASER3 produced the best result for baseline experiments, and its results consistently improved after applying heuristics.
@@
In Table 7121212Since for the SiTa pair the disparity observed was marginal, we exclude this language-pair from our discussion., the best result of each heuristic category reduces this disparity for all experiments related to CCAlign. Disparity has been reduced on average by 95.87% for CCAligned-XLMR/LaBSE and CCMatrix-LaBSE, compared to LASER3. The disparity between CCMatrix-XLMR EnSi/EnTa models, compared to LASER3 was reduced by 48% only, meaning that XLM-R ranked corpus contains noise which cannot be mitigated by the heuristics alone. However, our hypothesis holds in most cases, and it is safe to say that heuristic-based PDC mitigate the bias brought in by the multiPLM.
@@
Disparity (ChrF++) +/-Disparity wrt Baseline (%) Disparity (ChrF++) +/- Disparity wrt Baseline (%)
@@
Table 7: Disparity in ChrF++ points, among the NMT models (XLM-R/LaBSE) with the best scores with respect to LASER3 after applying the individual/combined heuristics. The +/- Disparity is this disparity as a percentage of the baseline disparity.

### v3

@@ 3 Methodology
Improved Taxonomy: Although translation pairs can have overlapping URLs, acronyms, etc, excessive inclusions of such content in a sentence (e.g. consider the sentence ‘Contact: Diane Anderson 076-8268914, info@sandnasbadenscamping.se’ [^2] More examples are in Table 11 in Appendix B.) do not provide meaningful content for an NMT system to learn from. However, under  Ranathunga et al. (2024)’s taxonomy, such sentence pairs would likely be categorised as perfect translation-pairs (CC). Therefore, we define a new noise category CCN (high-overlapping non-translatable text) to capture such sentence pairs.
Secondly, we consider the upper limit for short sentences as five words [^3] We considered thresholds 3,4 and 5 and empirically selected this threshold as it gave the highest NMT gains. Results in Table 12 in Appendix C.. Finally, we improve the definition of WL (wrong language) to consider a threshold in determining whether a sentence pair should be marked as wrong language. Table 3 shows these changes. The complete list of these noise categories and example parallel sentences are available in Table 9 and Table 10 (respectively) in Appendix B.
Selection of Heuristics:
Table 4 shows how the heuristics discussed in Section 2.3 may help in removing different noise categories. Note that a deduplication-based heuristic cannot be associated with any in the taxonomy, as it does not apply to individual sentence pairs. In addition to the deduplication strategies discussed in Section 2.3, we introduce an n-gram-based deduplication, meaning that sentences would be removed if they overlap in a consecutive n-gram text span.
@@
High-overlapping non-translatable text CCN LID, sentWRatio, sentCRatio
@@
We conduct a human evaluation to quantify the noise before and after applying the heuristics. The annotator selection criteria, resources and training provided, the payment details, etc are described in Appendix A.
For each language-pair, we obtain top 1000 samples in each of the ranked corpora using embeddings obtained from LASER3, XLM-R, and LaBSE, we randomly select 100 parallel sentences for each language pair. We ask the translators to annotate each sentence pair using the taxonomy discussed in Section 3.
Each sentence pair is annotated by three translators to reduce any potential bias inherent in the individual translators. For the three annotators, the Fleiss Kappa scores are 0.833 for EnSi, 0.651 for EnTa, and 0.649 for SiTa. We note that the results for EnTa and SiTa are very close.
@@ 4.1 Data
We use the language pairs, En-Si, En-Ta, and Si-Ta in our experiments. Sinhala and Tamil are morphologically rich, low-resource and mid-resourced languages Joshi et al. (2020); Ranathunga and de Silva (2022), respectively. Languages are selected considering the availability of human evaluators. We select CCMatrix and CCAligned as the web-mined corpora. Both these corpora include parallel data for the language pairs considered in the research. More details of these datasets and language pairs are shown in Appendix D.
For the NMT experiments, we use the dev and devtest subsets from the Flores-200 Costa-jussà et al. (2022) dataset [^4] https://github.com/openlanguagedata/flores as validation and evaluation sets, respectively. Dataset statistics are provided in Table 5.
@@
4.2 Selection of muliPLMs
We select LASER3, XLM-R, and LaBSE for obtaining embeddings for the sentences to determine the semantic similarity. XLM-R, different to others, was trained purely on monolingual data, but has proven to be useful for cross-lingual tasks as well Choi et al. (2021); Conneau et al. (2020). All three models include En, Si, and Ta. Details of the multiPLMs are shown in Appendix E.
4.3 Heuristic-based PDC Experiments
Each heuristic is applied independently to the source (S), target (T), and both sides (ST) of the corpora. In line with the original sentence alignment conducted for CCMatrix and CCAligned, we treat En as the source side. For Si-Ta, Si is considered the source because it is more common for a Si sentence to be translated to Ta Farhath et al. (2018). Finally, for each multiPLM, the retained sentences are ranked in descending order with cosine similarity.
Deduplication: We consider different granularities of deduplication. i.e. identical deduplication (dedup), deduplication after removing numbers only (nums) and removing both numbers and punctuations (punctsNums). Subsequently, we deduplicate considering different n-gram spans, i.e. 4-grams, 5-grams, 6-grams and 7-grams.
Length-based: We filter short sentences less than five words3. While some research has suggested removing extremely long sentences (Minh-Cong et al., 2023; Gala et al., 2023), we find that the percentage of longer sentences is lower and that removing them has a negligible effect. Thus, this result is not reported.
LID-based: We use a public LID model [^5] https://github.com/facebookresearch/fairseq/tree/nllb Costa-jussà et al. (2022) to predict the language of each sentence. The predicted label is then used as a standalone heuristic (LID) and in combination with its associated prediction probability (LIDThresh), with threshold of 0.7666Thresholds below 0.7 reduce NMT results..
Ratio-based: For STRatio, 0.79-1.39, 0.87-1.62 and 0.85-1.57 were selected as thresholds for En-Si, En-Ta and Si-Ta, respectively. These thresholds are determined by calculating the mean and the standard deviation obtained for the validation set in a human-crafted trilingual dataset (Fernando et al., 2020; Ranathunga et al., 2018). Following observations of Hangya and Fraser (2018), 0.6 is selected as the threshold for sentWRatio and  sentCRatio.
4.4 NMT Experiments
First, a Sentencepiece [^7] https://github.com/google/sentencepiece tokenizer with a vocabulary size of 25000 is trained. Then we use the fairseq toolkit  (Ott et al., 2019) to model and train transformer-based Seq-to-Seq NMT models until convergence. Hyperparameters used in the NMT experiments are shown in Table 13 in Appendix F. The baseline NMT models are trained on the top 100,000 sentence pairs from the ranked corpus. We use ChrF++ Popović (2017) to report NMT results.
5 Results and Analysis
We report the results obtained for the NMT models trained in the forward direction. The results of the experiments are in Table 6. Heuristic-wise best result is summarised in Table 14 in Appendix G. As evident from these tables, as well as from Figure 1, NMT results across different multiPLMs show a great disparity in the baseline NMT scores for En-Si and En-Ta language pairs. In the following sub-sections, we discuss how the use of heuristics is useful in mitigating this disparity and improving overall NMT results.
5.1 Impact of Heuristics on NMT Results
5.1.1 Impact of Deduplication-based PDC
We observe that deduplication, irrespective of the heuristic-applied side (S/T/ST), outperforms the baseline in 89% of the experiments. Overall, dedup considering both source and target seems to be the most effective — it outperforms the baseline in 94% of the experiments. In comparison, dedup target-only and source-only outperform the baseline in 89% and 83% of the experiments, respectively.
@@ 5.1.1 Impact of Deduplication-based PDC
We apply our newly introduced n-gram-based deduplication dedup+ngram, on top of  dedup. We find that for each result column, there is a dedup+ngram result that outperforms the best results obtained with the corresponding dedup result. To observe the impact of the n value on the NMT result, we plot Figure 2 showing the percentage of n-gram experiments exceeding the highest dedup result, with respect to the multiPLM (language-wise result is in Figure 4 of the Appendix G). We observe a consistent pattern - n=5 or 6 perform the best in a majority of cases. We believe that 4-gram results in an overly aggressive deduplication. However, the exact n-gram depends on the corpus characteristics.
We observe that dedup+punctNums outperforms  dedup+nums and dedup in 78% and 89% of the experiments (respectively), proving that (dedup+punctNums) to be more impactful. Finally, we analyse the impact of dedup+puntsNums+ngram888puntsNums and ngram has been applied on top of dedup.. Compared to other dedup combinations999dedup, dedup+ngram, nums, and dedup+punctNums, dedup+punctNums+ngram produce the best result across 67% of the experiments.
@@ 5.1.2 Impact of Length-Based PDC
sLength surpasses the baseline in 89% of the experiments. Therefore, we can conclude that sLength is favourable as a heuristic. When analysing the side on which the heuristic is applied, we observe that applying it to both ST is the most effective (similar to dedup), followed by T and then S (56%, 28%, and 17% of the experiments, respectively).
Recall that our manual inspection noticed that XLM-R and LaBSE tend to prioritise shorter sentences. This observation is affirmed by the sLength results. For En-Si and En-Ta, this heuristic resulted in substantial gains for many of the XLM-R and LaBSE experiments, while it shows marginal improvements for LASER3.
@@ 5.1.3 Impact of LID-Based PDC
With LID-based PDC, we observe substantial improvements for XLM-R and LaBSE, except for Si-Ta, for which the gains are marginal. Gains are reported by LASER3 as well, though not very significant. The highest gain of +20.61 ChrF++ is reported for XLM-R for the CCAligned-EnTa corpus, while a gain of +11.40 is reported by LaBSE for the same corpus.
NMT models trained after applying  LID outperform the baseline in 85% of the experiments, while  LIDThresh outperforms LID in 72% of the experiments. Therefore, we conclude that LIDThresh would be the most suitable heuristic. We observe that the gains for LIDThresh with respect to LID are least with Si-Ta with 50% while for En-Si and En-Ta it is 83% each. We assume that this is due to a limitation with the LID model, which is not optimised for Si and Ta.
@@ 1. Impact of the Individual Heuristics on the NMT Results:
The LID-based heuristic emerged as the most impactful in 44% of the experiments. Deduplication and sentence-length heuristics were the most effective in 33% and 17% of the experiments, respectively. The ratio-based heuristic alone did not hold superior results compared to others, making it the least impactful single heuristic. There is no single heuristic that consistently produces the best gains for all scenarios. For LASER3, XLM-R, and LaBSE, we observe the highest ChrF++ gains of +10.57, +20.61, and +11.40 for CCMatrix-EnTa and CCAligned-EnTa and CCAligned-EnTa corpora, respectively.
2. Impact of Combined Heuristics:
@@
The combination of heuristics produced the best score, compared to the best-performing individual heuristics, except for the CCMatrix-SiTa. In this, the combination reduced the dataset size by 54% (Table 15 in Appendix G), resulting in only 98k parallel sentences. We suspect this reduction in dataset size and the residual noise could result in the reduced score.
@@
CCMatrix-SiTa gains across multiPLMs are marginal. The same pattern holds with LASER3 across En-Ta and En-Si languages, with the exception of CCMatrix-EnTa corpus. Since LASER3 is already trained with OPUS data Tiedemann and Thottingal (2020) [^10] https://opus.nlpl.eu/, it is believed that LASER3 is already optimised Moon et al. (2023) towards ranking the CCMatrix and CCAligned corpora better. This is further affirmed with the relatively higher NMT scores of the baseline experiments.
For the rest of the ranked corpora, the best NMT scores were reported when the combined heuristic was applied. It was noted the combination always had heuristics, dedup+punctNums+(n)gram+sLength+LIDThresh as a common combination for 77% of the experiments. The specific n-value was dependent on the dataset. In the combination producing the best gains, the ratio-based heuristic was different based on the dataset/language pair. In 61% of the cases, the best scores were obtained with sentWRatio, while STRatio yielded the best results in 16% of the experiments.
@@
The combination performed best without the LID heuristic only in CCAligned-SiTa with XLM-R. However, we include LIDThresh into the combination for two reasons. (1) Even with LIDThreshold, the score only lags by (-0.19) ChrF++ scores compared to the best NMT score, which is negligible. (2) The most influential individual heuristic is LIDThresh-based for most cases.
Therefore we recommend the heuristic combination, dedup+punctNums+(n)gram+sLength
+LIDThresh+sentWRatio
to be applied as the rule-based heuristic combination for the PDC task.
3. Reducing the Disparity Across multiPLMs:
We calculate the disparity ($\Delta$) as the difference in NMT scores between LASER3 and the XLM-R or LaBSE scores. Equation 1 shows the baseline disparity calculation, where $BL_{\text{{multiPLM}}}$ refers to the baseline NMT score obtained using embeddings from either XLM-R or LABSE during ranking.
$\Delta_{\textit{Baseline}}=BL_{\text{{LASER3}}}-BL_{\text{{multiPLM}}}$ (1)
The disparity for each heuristic is calculated with respect to LASER3 and the best NMT score obtained from either XLM-R or LaBSE. The results are shown in Table 7. Additionally, we calculate the disparity reduction percentage ($\Delta~Reduction(\%)$) as defined in Equation 2. Here, $\Delta_{heuristic}$ is the disparity corresponding to the best-performing NMT models after applying the respective heuristic or the optimal combination.
$\Delta~Reduction(\%)=\frac{\Delta_{\text{baseline}}-\Delta_{\text{heuristic}}}{\Delta_{\text{baseline}}}\times 100\%$ (2)
@@ 3. Reducing the Disparity Across multiPLMs:
In Table 7111111Since the disparity observed was marginal for the SiTa pair, we exclude this language pair from our discussion., we observe that the disparity has been reduced on average by 95.87% for CCAligned-XLM-R/LaBSE and CCMatrix-LaBSE, compared to LASER3. The disparity between CCMatrix-XLM-R EnSi/EnTa models, compared to LASER3 was reduced by only 48%, meaning that XLM-R ranked corpus contains noise that cannot be mitigated by the heuristics alone. However, our hypothesis holds in most cases, and it is safe to say that heuristic-based PDC mitigates the bias brought in by the multiPLM.
@@
$\Delta$ (ChrF++) $\Delta$ Reduction (%) $\Delta$ (ChrF++) $\Delta$ Reduction (%)
@@
Table 7: Disparity ($\Delta$) in ChrF++ points, among the NMT models (XLM-R/LaBSE) with the best scores with respect to LASER3 after applying the individual/combined heuristics. The $\Delta$ Reduction (%) is this disparity as a percentage of the baseline disparity.